# Binary superfluids: low-energy properties and dissipative processes from spontaneous emission of massive phonons

### v1

@@ VI Spontaneous phonon emission
Having obtained the low-energy effective theory for binary superfluids, we now consider the effect of a background velocity profile. In particular, we are interested to determine the effect of the phonon mass on the spontaneous emission process at the acoustic horizon Liberati et al. (2006a); Weinfurtner et al. (2007). Given the two Lagrangians in Eqs. (124) and (125) for the massless and massive phonons, respectively, the corresponding acoustic horizons are in the positions, $x^{H}_{a}$ and $x^{H}_{b}$, such that
@@
where we assumed that the flow is along the $x-$direction and thus the acoustic horizons are two planes centered at $x^{H}_{a}$ and $x^{H}_{b}$. More precisely, the horizon positions can be determined with precision of the order of the healing length, that is the threshold scale length required for the validity of the hydrodynamic description of realistic superfluids Dalfovo et al. (1999). Close to the acoustic horizons, we consider that the fluid velocity is given by $\bm{v}=-v\bm{\hat{}}x$ with velocity profile
@@ VI.1 Dispersion close to the acoustic horizon
Here we present a qualitative description of the phenomenon, based on semi-classical arguments. For this purpose, it is useful to analyze first the dispersion law of the phonons around the acoustic horizon. The equation of motion of phonons with mass $m_{\text{ph}}$ is
@@
which is valid in the collinear regime, meaning that both the fluid velocity and the phonon momentum are along the $x$–axis. In the following, we restrict ourselves, to positive energy states in the nonrelativistic limit ($v\ll 1$ and $c_{s}\ll 1$). In this case the dispersion law of phonons simplifies:
@@
which has still a relativistic-like form, because we did not assume that $|p_{x}|<m_{\text{ph}}c_{s}$. In this way we can take into account the case of an arbitrarily small phonon mass. The corresponding group velocity is given by
@@
where in the last equation we exploited Eq.  (134) ($x$ being the distance form the horizon). Clearly, for $m_{\text{ph}}=0$ we have $p_{\text{c}}=0$ and this is independent of $x$. Instead, for nonvanishing phonon mass, the above expression implies that the largest threshold momentum is obtained at the shorter distance from the horizon, that is for $x\simeq L_{c}$, where $L_{c}$ is a distance, already introduced after Eq. (134), of the order of the healing length Dalfovo et al. (1999) (see also Son and Stephanov (2002a) for a discussion of healing lengths in binary superfluids with Rabi coupling).
From the relation in Eq. (140), it follows that the threshold energy is
$E_{c}=-p_{c}v+\sqrt{p_{c}^{2}c_{s}^{2}+m_{\text{ph}}^{2}c_{s}^{4}}\simeq m_{\text{ph}}c_{s}^{3/2}\sqrt{4\pi TL_{c}}\,,$ (141)
where we have considered the regime
$c_{s}\ll\frac{1}{4\pi TL_{c}}\,,$ (142)
so that the phonon threshold momentum is much larger than $m_{\text{ph}}\,c_{s}$. In other words, the phonon mass produces a small variation of the phonon energy.
@@ VI.2 Thermodynamics close to the horizon
takes into account that phonons are only emitted in the direction orthogonal to the horizon Mannarelli et al. (2021b); L. Chiofalo et al. (2024), that is along the $x$–axis, and that the phonon momenta must be larger than the threshold momentum, given in Eq. (140). Upon changing variable, $p_{x}\to E$, and then defining $z=E/T$, this expression can be rewritten as
@@
$\frac{3E_{c}}{\pi^{2}T}=\frac{6m_{\text{ph}}c_{s}}{\pi^{3/2}}\sqrt{\frac{c_{s}L_{c}}{T}}\,,$ (148)
then ruled not only by the Rabi coupling, entering the phonon mass, but also by the ratio between $L_{c}$ and the Hawking temperature. Note that the correction is nonperturbative, in the sense that it is of order $\lambda^{1/2}$, see Eqs. (126) and (141). This result is similar to the one found in the calculation of the free-energy of a scalar field in thermal field theory, see for instance Le Bellac (2000). Such behavior reflects the breakdown of perturbation theory due to the presence of infrared divergences.
We can estimate the energy density reduction due to the Rabi coupling assuming that $L_{c}=\xi$, where $\xi$ is the healing length. In the set-up of Muñoz de Nova et al. (2019), where the experiment is performed with 87Rb atoms, the healing length is $\xi\simeq 2\,\mu$m, the speed of sound is $c_{s}\simeq 0.3$ mm/s, and the estimated Hawking temperature is $T\simeq 0.1\,mc_{s}^{2}\simeq 0.1$ nK, where $m$ is the mass of a 87Rb atom. Then, considering that $m_{\text{ph}}c_{s}\simeq 2\sqrt{\lambda}$ and that $\lambda=m\tilde{\lambda}$ (see Eq. (34)) with $\tilde{\lambda}\sim 10^{-16}\,\text{eV}$ (see Eq. (35)), we have that
$\frac{3E_{c}}{\pi^{2}T}\simeq 0.2\,,$ (149)
therefore the effect is a reduction of the order of $20\%$ with respect to the same value for vanishing Rabi coupling. Note that the effect is proportional to $\tilde{\lambda}^{1/2}$ and it is independent of the mass of the used atoms. In a similar way, we can expand the phonon pressure
@@ VII Viscosity and the conformal limit close to the horizon
As a consequence of the spontaneous phonon emission close to the acoustic horizon, the transonic superfluid loses energy because a non-equilibrium (relaxation) dynamics sets in Mannarelli et al. (2021a, b); L. Chiofalo et al. (2024): the superfluid kinetic energy is transformed in heat. At the effective level, this behavior can be characterized by effective bulk, $\zeta_{\text{eff}}$, and shear, $\eta_{\text{eff}}$, viscous coefficients, associated to the spontaneous phonon emission. As shown in  L. Chiofalo et al. (2024), the shear-to-entropy density ratio, $\eta_{\text{eff}}/s_{\text{ph}}$, of the phonon gas emitted at the acoustic horizon saturates the so-called KSS bound. First, we briefly recap the results of  L. Chiofalo et al. (2024), obtained for a single superfluid and then move to the evaluation of the ratios $\zeta_{\text{eff}}/s_{\text{ph}}$ and $\eta_{\text{eff}}/s_{\text{ph}}$ in binary superfluids.
@@
where we have taken into account that phonons are mostly emitted along the direction orthogonal to the horizon and on the r.h.s. we have used Eqs. (131) and (134). Equating $\sigma_{xx}^{\prime}=P_{\text{ph}}$, where $P_{\text{ph}}$ is given in Eq. (150), we have that
@@
which relates the bulk viscosity to the pressure of the phonon gas. This equation states that close to the acoustic horizon the only considered dissipative mechanism is the spontaneous phonon emission; any other viscous process is assumed negligible. However, it is well known that scattering processes are dissipative and contribute to the fluid viscosity Khalatnikov (1965). Here we consider the phonon gas in a region close to the acoustic horizon of extension smaller that the typical phonon mean free path so that phonon scattering processes can be neglected. Then, we use the thermodynamic relation
@@
$2\pi T\zeta_{\text{eff}}=\frac{Ts_{\text{ph}}}{2}\,.$ (158)
@@
In a similar way, we can obtain the ratio between the shear viscosity coefficient and the entropy density. In this case, we have to assume that the flow has shear, thus that the velocity, which is along the $x$–direction, depends on a transverse coordinate, for instance $v_{x}$ depends on the $y$–coordinate. In this conditions, there is a ”transverse” pressure L. Chiofalo et al. (2024), such that
@@
For nonvanishing Rabi coupling, one of the two phonons becomes massive and this is expected to change the viscosity to entropy density ratios. The reason is that for massive phonons the relation in Eq. (152) does not hold. The basic reason is that the phonon mass breaks the scale invariance. More in details, the needed ingredients to saturate the KSS bound in the presence of an acoustic horizon are the followings:
@@
While the conditions i) and ii) hold as well for a massive phonon gas, the condition iii) is violated by the phonon mass. As we have shown in the previous section, a mass term implies Eq. (153), hence
@@
and using the estimate in Eq. (149), we have a reduction of about $10\%$ with respect to the KSS bound. This is the main result of the present work.
Let us remark an important aspect of the obtained result: it holds because the phonon mass is not an interaction term, therefore it does not produce phonon scatterings. Its only effect is to break the scale invariance and thus the phonon pressure must be less than the phonon energy density. This means that the bulk and shear viscosities can still be written as in Eqs. (158) and  (160), respectively, but now the phonon pressure is given by Eq. (151). Since the phonon mass reduces the phonon pressure, it follows that with increasing phonon mass the effective viscosities are reduced, as well.
Concerning the effect of the interactions on the KSS bound, we observe that they immediately violate the conditions i) and iii) (encoding scale invariance), therefore they are expected to induce deviations from the same bound. This aspect deserves future investigation.
@@ VIII Conclusions
Binary ultracold gases are versatile playgrounds for producing nontrivial phases of matter  Yago Malo et al. (2024). In the present paper, we have determined the properties of the superfluid ground-state for a two-boson mixture coupled by intra- and inter-species interactions, as well as by a Rabi exchange term. Later on, the corresponding effective low-energy excitation theory has been determined. Assuming a perturbative Rabi coupling, we have shown that when the number densities of the two fluids are similar, it is possible to write the total pressure of the system as the sum of two partial pressures. Moreover, we have obtained that in the same conditions, the low-energy excitations of the binary superfluid are two decoupled modes, corresponding to the relative and the total phase oscillations. The former is massive, with a squared mass proportional to the Rabi coupling, while the latter remains massless.
Separating the total system as the sum of two noninteracting subsystems is useful when discussing the spontaneous phonon emission close to an acoustic horizon. Indeed, we can treat the emission of the two phonon modes as independent processes. The spontaneous phonon emission is an irreversible process, that at the hydrodynamic level amounts to effective viscosities. In binary superfluids, the spontaneous emission of massless phonons is analogous to the phonon process in a single superfluid, thus the viscosity-to-entropy ratios saturates the KSS bound Kovtun et al. (2005); L. Chiofalo et al. (2024): it is equal to $1/4\pi$. However, the spontaneous emission of massive phonons at the acoustic horizon violates the same bound: in this case, ${\eta_{\rm eff}}/{s_{\rm ph}}={\zeta_{\rm eff}}/{s_{\rm ph}}<1/4\pi$. The violation is estimated of the order of $10\%$ using the specific setup in  Muñoz de Nova et al. (2019). In general, the magnitude of the same deviation scales with the square root of the Rabi frequency, and it is independent on the mass of the used atoms. Therefore, it seems that to enhance the deviation from the KSS bound one needs to design systems with the highest possible Rabi frequency and with sufficiently tunable speeds of sound and/or fluid velocity as necessary ingredients to have an acoustic horizon. Our setup provides a versatile tool to explore the degree of KSS-bound violation and even advance the exciting quest for a universal $\eta/s$ bound, under controllable external conditions that can be engineered in current quantum technology platforms.
@@
Acknowledgments – The authors are pleased to thank Stefano Carretta, Leonardo Fallani, Stefano Liberati, Luca Salasnich, Paolo Santini, and Andrea Trombettoni for support and fruitful discussions.
@@
This study has been also carried out within the National Centre on HPC, Big Data and Quantum Computing - SPOKE 10 (Quantum Computing) and received funding from the European Union Next-GenerationEU - National Recovery and Resilience Plan (NRRP) – MISSION 4 COMPONENT 2, INVESTMENT N. 1.4 – CUP N. I53C22000690001.

### v2

@@ VI Spontaneous phonon emission
Having obtained the low-energy effective theory for binary superfluids, we consider the effect of a background velocity profile. In particular, we are interested in determining the effect of the phonon mass on the spontaneous emission process at the acoustic horizon Liberati et al. (2006a); Weinfurtner et al. (2007). Given the two Lagrangians in Eqs. (126) and (127) for the massless and massive phonons, respectively, the corresponding acoustic horizons are in the positions, $x^{H}_{a}$ and $x^{H}_{b}$, such that
@@
where we assumed that the flow is along the $x$ direction and thus the acoustic horizons are two planes centered at $x^{H}_{a}$ and $x^{H}_{b}$. More precisely, the horizon positions can be determined with precision of the order of the healing length, that is the threshold scale length required for the validity of the hydrodynamic description of realistic superfluids Dalfovo et al. (1999). Close to the acoustic horizons, we consider that the fluid velocity is given by $\bm{v}=-v\bm{\hat{}}x$ with velocity profile
@@ VI.1 Dispersion close to the acoustic horizon
Here we present a qualitative description of the phenomenon, based on semiclassical arguments. For this purpose, it is useful to analyze first the dispersion law of the phonons around the acoustic horizon. The equation of motion of phonons with mass $m_{\text{ph}}$ is
@@
which is valid in the collinear regime, meaning that both the fluid velocity and the phonon momentum are along the $x$ axis. In the following, we restrict ourselves, to positive energy states in the nonrelativistic limit ($v\ll 1$ and $c_{s}\ll 1$). In this case the dispersion law of phonons simplifies:
@@
which has still a relativisticlike form, because we did not assume that $|p_{x}|<m_{\text{ph}}c_{s}$. In this way we can take into account the case of an arbitrarily small phonon mass. The corresponding group velocity is given by
@@
where in the last equation we exploited Eq.  (136) ($x$ being the distance from the horizon). Clearly, for $m_{\text{ph}}=0$ we have $p_{\text{c}}=0$, and this is independent of $x$. Instead, for nonvanishing phonon mass, the above expression implies that the largest threshold momentum is obtained at the shorter distance from the horizon, that is for $x\simeq L_{c}$, where $L_{c}$ is a distance, already introduced after Eq. (136), of the order of the healing length Dalfovo et al. (1999) (see also Son and Stephanov (2002a) for a discussion of healing lengths in binary superfluids with Rabi coupling). The threshold momentum $p_{c}$ in Eq. (142) must be lower than the inverse of the healing length,
$p_{c}\ll\frac{1}{\xi}\,,$ (143)
to ensure the validity of the hydrodynamic description. In typical superfluid experiments $T\sim 0.1\,mc_{s}^{2}$teinhauer (2016); Muñoz de Nova et al. (2019), and using Eq. (142) we have that the above inequality can be rewritten as
$m_{\text{ph}}c_{s}^{2}\ll 10\,T\,,$ (144)
which is satisfied because $m_{\text{ph}}c_{s}^{2}\lesssim 2\sqrt{m\tilde{\lambda}}c_{s}<T$, where we used $m\simeq 100$ GeV and $\tilde{\lambda}\simeq 10^{-16}$ eV.
Upon substituting Eq. (142) in Eq.(140), it follows that the threshold energy is
$E_{c}=$ $-p_{c}v+\sqrt{p_{c}^{2}c_{s}^{2}+m_{\text{ph}}^{2}c_{s}^{4}}\,\simeq$
$\simeq m_{\text{ph}}c_{s}^{2}\left(\frac{\sqrt{k}}{2}+\sqrt{1+\frac{1}{k}}-\frac{1}{\sqrt{k}}\right)\,,$ (145)
where
$k=\frac{4\pi\,T\,L_{c}}{c_{s}}\,.$ (146)
Therefore the threshold energy is ruled not only by the Rabi coupling, entering in the phonon mass, but also by $L_{c}$, the Hawking temperature, as well as by the speed of sound $c_{s}$.
@@ VI.2 Thermodynamics close to the horizon
takes into account that phonons are only emitted in the direction orthogonal to the horizon Mannarelli et al. (2021b); L. Chiofalo et al. (2024), that is along the $x$ axis, and that the phonon momenta must be larger than the threshold momentum, given in Eq. (142). Upon changing the variable, $p_{x}\to E$, and then defining $z=E/T$, this expression can be rewritten as
@@
$\frac{3E_{c}}{\pi^{2}T}\simeq\frac{3m_{\text{ph}}c_{s}^{2}}{T\pi^{2}}\left(\frac{\sqrt{k}}{2}+\sqrt{1+\frac{1}{k}}-\frac{1}{\sqrt{k}}\right)\,,$ (152)
where $k$ is the quantity defined in Eq. (146), thus it depends on $T$ as well as on $L_{c}$ and $c_{s}$. Note that the correction is nonperturbative, in the sense that it is of order $\lambda^{1/2}$, see Eqs. (128) and (VI.1). This result is similar to the one found in the calculation of the free energy of a scalar field in thermal field theory, see for instance Le Bellac (2000). Such behavior reflects the breakdown of perturbation theory due to the presence of infrared divergences.
We can estimate the energy density reduction due to the Rabi coupling assuming that $L_{c}=\xi$, where $\xi$ is the healing length. In the setup of Steinhauer (2016); Muñoz de Nova et al. (2019), where the experiment is performed with 87Rb atoms, the healing length is $\xi\simeq 2\,\mu$m, the speed of sound is $c_{s}\simeq 0.3$ mm/s, and the estimated Hawking temperature is $T\simeq 0.1\,mc_{s}^{2}\simeq 0.1$ nK, where $m$ is the mass of a 87Rb atom. Upon substituting these values in Eq. (146), we obtain that $k\simeq 0.9$. Then, considering that $m_{\text{ph}}c_{s}^{2}<T$ we find that
$\frac{3E_{c}}{\pi^{2}T}\lesssim 0.2\,.$ (153)
Therefore, the effect is a reduction of the order of $20\%$ with respect to the same value for vanishing Rabi coupling. Note that the effect is actually proportional to $\tilde{\lambda}^{1/2}$ and it is independent of the mass of the used atoms. In a similar way, we can expand the phonon pressure
@@ VII Viscosity and the conformal limit close to the horizon
As a consequence of the spontaneous phonon emission close to the acoustic horizon, the transonic superfluid loses energy because a nonequilibrium (relaxation) dynamics sets in Mannarelli et al. (2021a, b); L. Chiofalo et al. (2024): the superfluid kinetic energy is transformed in heat. At the effective level, this behavior can be characterized by effective bulk, $\zeta_{\text{eff}}$, and shear, $\eta_{\text{eff}}$, viscous coefficients, associated to the spontaneous phonon emission. As shown in  L. Chiofalo et al. (2024), the shear-to-entropy density ratio, $\eta_{\text{eff}}/s_{\text{ph}}$, of the phonon gas emitted at the acoustic horizon saturates the so-called KSS bound. First, we briefly recap the results of  L. Chiofalo et al. (2024), obtained for a single superfluid and then move to the evaluation of the ratios $\zeta_{\text{eff}}/s_{\text{ph}}$ and $\eta_{\text{eff}}/s_{\text{ph}}$ in binary superfluids.
@@
where we have taken into account that phonons are mostly emitted along the direction orthogonal to the horizon and on the rhs we have used Eqs. (133) and (136). Equating $\sigma_{xx}^{\prime}=P_{\text{ph}}$, where $P_{\text{ph}}$ is given in Eq. (154), we have that
@@
which relates the bulk viscosity to the pressure of the phonon gas. This equation states that close to the acoustic horizon the only considered dissipative mechanism is the spontaneous phonon emission; any other viscous process is assumed negligible. However, it is well known that scattering processes are dissipative and contribute to the fluid viscosity Khalatnikov (1965). Here we consider the phonon gas in a region very close to the acoustic horizon. The extension of this region is assumed to be smaller than the typical phonon mean free path, so that phonon scattering processes can be neglected. Then, we use the thermodynamic relation
@@
$2\pi T\zeta_{\text{eff}}=\frac{Ts_{\text{ph}}}{2}\,,$ (162)
@@
In a similar way, we can obtain the ratio between the shear viscosity coefficient and the entropy density. In this case, we have to assume that the flow has shear, thus that the velocity, which is along the $x$ direction, depends on a transverse coordinate, for instance $v_{x}$ depends on the $y$ coordinate. In this condition, there is a ”transverse” pressure L. Chiofalo et al. (2024), such that
@@
For nonvanishing Rabi coupling, one of the two phonons becomes massive and this is expected to change the viscosity to entropy density ratios. The reason is that for massive phonons the relation in Eq. (156) does not hold. The basic reason is that the phonon mass breaks the scale invariance. More specifically, the needed ingredients to saturate the KSS bound in the presence of an acoustic horizon are the following:
@@
While conditions i) and ii) hold as well for a massive phonon gas, condition iii) is violated by the phonon mass. As we have shown in the previous section, a mass term implies Eq. (157); hence
@@
and using the estimate in Eq. (153), we have a reduction of about $10\%$ with respect to the KSS bound. This is the main result of the present work. It has been obtained assuming that the two horizons are sufficiently distant, so that one can separately account for their contributions to entropy and shear viscosity. In the case where the two horizons are very close or coincide, one should consider the combined effect of the emission processes of massless and massive phonons. Since the two emission processes are independent, the total entropy and shear viscosity are given by the sum of the corresponding contributions. This leads to a factor $2$ reduction of the KSS bound violation reported in Eq. (167).
Let us remark on an important aspect of the obtained results: they hold because the phonon mass is not an interaction term; therefore it does not produce phonon scatterings. Its only effect is to break the scale invariance and thus the phonon pressure must be less than the phonon energy density. This means that the bulk and shear viscosities can still be written as in Eqs. (162) and  (164), respectively, but now the phonon pressure is given by Eq. (155). Since the phonon mass reduces the phonon pressure, it follows that with increasing phonon mass the effective viscosities are reduced, as well.
Concerning the effect of the interactions on the KSS bound, we observe that they immediately violate the conditions i) and iii) (encoding scale invariance); therefore they are expected to induce deviations from the same bound. This aspect deserves future investigation.
@@ VIII Conclusions
Binary ultracold gases are versatile playgrounds for producing nontrivial phases of matter Yago Malo et al. (2024). In the present paper, we have determined the properties of the superfluid ground state for a two-boson mixture coupled by intraspecies and interspecies interactions, as well as by a Rabi exchange term. Later on, the corresponding effective low-energy excitation theory was determined. Assuming a perturbative Rabi coupling, we have shown that when the number densities of the two fluids are similar, it is possible to write the total pressure of the system as the sum of two partial pressures. Moreover, we have obtained that in the same conditions, the low-energy excitations of the binary superfluid are two decoupled modes, corresponding to the relative and the total phase oscillations. The former is massive, with a squared mass proportional to the Rabi coupling, while the latter remains massless.
Separating the total system as the sum of two noninteracting subsystems is useful when discussing the spontaneous phonon emission close to an acoustic horizon. Indeed, we can treat the emission of the two phonon modes as independent processes. The spontaneous phonon emission is an irreversible process, that at the hydrodynamic level amounts to effective viscosities. In binary superfluids, the spontaneous emission of massless phonons is analogous to the phonon process in a single superfluid, thus the viscosity-to-entropy ratios saturate the KSS bound Kovtun et al. (2005); L. Chiofalo et al. (2024): it is equal to $1/4\pi$. However, the spontaneous emission of massive phonons at the acoustic horizon violates the same bound: in this case, ${\eta_{\rm eff}}/{s_{\rm ph}}={\zeta_{\rm eff}}/{s_{\rm ph}}<1/4\pi$. The violation is estimated of the order of $10\%$ using the specific setup in  Muñoz de Nova et al. (2019). In general, the magnitude of the same deviation scales with the square root of the Rabi frequency, and it is independent on the mass of the used atoms. Therefore, it seems that to enhance the deviation from the KSS bound one needs to design systems with the highest possible Rabi frequency and with sufficiently tunable speeds of sound and/or fluid velocity as necessary ingredients to have an acoustic horizon. Our setup provides a versatile tool to explore the degree of KSS-bound violation and even advance the exciting quest for a universal $\eta/s$ bound, under controllable external conditions that can be engineered in current quantum technology platforms.
@@
Acknowledgments
The authors are pleased to thank Stefano Carretta, Leonardo Fallani, Stefano Liberati, Luca Salasnich, Paolo Santini, and Andrea Trombettoni for support and fruitful discussions.
@@ Acknowledgments
This study has been also carried out within the National Centre on HPC, Big Data and Quantum Computing - SPOKE 10 (Quantum Computing) and received funding from the European Union Next-GenerationEU - National Recovery and Resilience Plan (NRRP) – MISSION 4 COMPONENT 2, INVESTMENT N. 1.4 – CUP N. I53C22000690001. M.L.C. acknowledges the Terrestrial Very-Long-Baseline Atom Interferometry (TVLBAI) proto-collaboration
Data availability
No data were created or analyzed in this study.